# Euclid Quick Data Release (Q1)

### v1

@@
The SPE processing function (PF) of the Euclid pipeline is dedicated to the automatic analysis of one-dimensional spectra to determine redshifts, line fluxes, and spectral classifications. The first Euclid Quick Data Release (Q1) delivers these measurements for all $H_{\scriptscriptstyle\rm E}<22.5$ objects identified in the photometric survey. In this paper, we present an overview of the SPE PF algorithm and assess its performance by comparing its results with high-quality spectroscopic redshifts from the Dark Energy Spectroscopic Instrument (DESI) survey in the Euclid Deep Field North. Our findings highlight remarkable accuracy in successful redshift measurements, with a bias of less than $3\times 10^{-5}$ in $(z_{\rm SPE}-z_{\rm DESI})/(1+z_{\rm DESI})$ and a high precision of approximately $10^{-3}$. The majority of spectra have only a single spectral feature or none at all. To avoid spurious detections, where noise features are misinterpreted as lines or lines are misidentified, it is therefore essential to apply well-defined criteria on quantities such as the redshift probability or the $\mathrm{H}\,\alpha$ flux and signal-to-noise ratio. Using a well-tuned quality selection, we achieve an 89% redshift success rate in the target redshift range for cosmology ($1.80.9<z<$ 1.8), which is well covered by DESI for $z<1.6$. Outside this range where the $\mathrm{H}\,\alpha$ line is observable, redshift measurements are less reliable, except for sources showing specific spectral features (e.g., two bright lines or strong continuum). The classification based on the spectroscopy alone is effective for galaxies (about 80% success rate), while it is currently less efficient for stars and quasars (< ${<}\,60$%). Ongoing refinements along the entire chain of PFs are expected to enhance both the redshift measurements and the spectral classification, allowing us to define the large and reliable sample required for cosmological analyses. Overall, the Q1 SPE results are promising, demonstrating encouraging potential for cosmology.
@@ 1 Introduction
One of the two main objectives of the Euclid mission (Euclid Collaboration: Mellier et al. 2024) is to measure galaxy clustering in the redshift interval $1.880.84<z<$ 1.88 (hereafter, the ‘target range’), using spectroscopic redshifts measured on the spectra provided by the Near Infrared Spectro-Photometer (NISP) instrument (Euclid Collaboration: Jahnke et al. 2024). The spectra have a resolution $R\simeq 500$, cover the wavelength range 1206–1892 $\mathrm{nm}$, and redshifts in the target range are measured by identifying the $\mathrm{H}\,\alpha$ emission line. NISP is a slitless spectrometer, which means that the spectra of all objects on the sky potentially leave a trace on the spectroscopic exposures. The only condition for a spectrum to be extracted is that a corresponding object is identified in the photometric catalogue constructed from Euclid imaging. This paper presents spectroscopic measurements and redshifts released in the framework of the first ‘Quick Release’ of Euclid products (Euclid Quick Release Q1 2025; Euclid Collaboration: Aussel et al. 2025), for objects brighter than $H_{\scriptscriptstyle\rm E}=22.5$ and observed over the area of the Euclid Deep Fields (EDSs) in the configuration of the Euclid Wide Survey (EWS). However, given the specific objectives of the mission, in terms of redshift range and emission-line flux, as well as wavelength domain and intrinsic quality of the spectra, these data cannot be used as a homogeneous set for blind statistical studies, and special attention must be paid to redshift values outside the target range or with $\mathrm{H}\,\alpha$ lines fainter than the nominal limit of the survey. In addition, we describe the output and performance of the sole spectroscopy channel, while the final analysis of Euclid will benefit from the combination with the high-quality visible and infrared images provided by both the VIS (Euclid Collaboration: Cropper et al. 2024) and NISP instruments.
The paper is organised as follows. Section 2 describes the various spectroscopic products that are present in the Euclid scientific archive, together with the main parameters that are useful in extracting the best measurements, as well as a brief description of the methods of redshift calculation. Section 3 gives results on the efficiency and accuracy of the redshift measurements, while the results of the spectroscopic classification are presented in Sect. 4. Section 5 gives information about the quality and performance of the emission-line fluxes, and Sect. 6 discusses the quality of the redshift measurement for various subsamples with respect to existing redshift catalogues.
@@ 2.1 SPE PF spectra handling and basic description
The Euclid data processing is handled by the Science Ground Segment, which is split into Processing Functions (PFs), each taking care of a specific aspect. The Spectroscopy PF (SPE PF) is in charge of redshift estimation, measurement-reliability calculation, and spectroscopic classification of all objects present in the main catalogue created by the ‘merge’ PF (MER PF, Euclid Collaboration: Romelli et al. 2025). SPE PF first reads the 1D spectra produced by SIR PF (Euclid Collaboration: Copin et al. 2025) by combining the four individual exposures taken with different orientations of the grisms (see Euclid Collaboration: Scaramella et al. 2022 for details on the observing strategy). The data provided by SIR PF also include flags describing the quality of each pixel in the spectrum and the number of individual exposures that actually contributed to its final value. For this Q1 release, only the spectra of objects with $H_{\scriptscriptstyle\rm E}\leq 22.5$ were extracted. At fainter magnitudes, the catalogues derived from NISP imaging suffer from significant contamination by fake sources, due to the persistence effect in infrared detectors. Spurious residuals from preceding spectroscopic exposures are detected on the following photometric images, thus preventing SIR PF from reliably extracting source positions at $H_{\scriptscriptstyle\rm E}>22.5$ (see Euclid Collaboration: Polenta et al. 2025 for a description of the NIR PF and the persistence phenomenon). Before proceeding with the redshift estimate, pixels that were flagged as “do not use” by SIR PF are discarded (see Euclid Collaboration: Copin et al. 2025 for details on flags). Following tests performed on simulated data and early spectra to eliminate as many artificial features as possible, only pixels that are built from at least three exposures are retained. This is the minimum number to guarantee a proper performance of the PF-SIR sigma-clipping combination.
@@
We also discard pixels that have the sixth bit of the quality flag activated: this corresponds to pixels at the edges of the spectrum, for which the absolute flux calibration curve has values close to zero, making calibration uncertain. Because of the current SIR configuration, this flag limits the available redshift interval to 0.9–1.8 in this release. Accounting also for the various edge effects (size of the fields, focal plane geometry, and gaps between detectors – see Fig. 1), yields a sample of 3.8 million processed spectra, among the 5.1 million objects in the $H_{\scriptscriptstyle\rm E}\leq 22.5$ catalogue, over the 63.1 deg2 Q1 area. This includes all spectra with at least one valid pixel. To approach a science-grade sample, a lower threshold to the number of valid pixels is clearly required; however, at this stage we chose not to apply any such cut, since future improvements in the pipeline may lead to different values for the optimal selection to be used. The nominal length of a spectrum that would not contain any invalid pixel is about 470 pixels.
@@
It is important to note that this sample of processed (and delivered) spectra is in fact nearly two orders of magnitude larger than the number of $\mathrm{H}\,\alpha$ emitters whose redshifts are expected to be measurable at the depth of the EWS, within the Q1 area. Following the original science requirements of the mission, the NISP spectrograph was designed as to secure a redshift for an average number of $1700$ galaxies per deg2, at a flux limit $f(\mathrm{H}\,\alpha)>2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$. Because of measuring redshifts using only the $\mathrm{H}\,\alpha$ line detection (see discussion in the following sections), by construction these galaxies will have a redshift within the target range $1.80.9<z<$ 1.8. In practice, this implies that within the Q1 release we would expect at best to recover a total of 63×1700100 000similar-to-or-equals63170010000063$\times$ 1700$\simeq$ 100\,00063 × 1700 ≃ 100 000 redshifts within the target range, i.e., about 2% of the total number of spectra extracted and delivered in Q1. The expected number of $\mathrm{H}\,\alpha$ emitters, in fact, should be even smaller, if we consider that the requirements are based on extracting all sources with broad-band magnitude $H_{\scriptscriptstyle\rm E}<24$, i.e., 1.5 mag deeper than the current selection. Thus, any scientific investigation based on these data requires a strong and careful selection, as we shall discuss in the following sections.
@@
To allow this selection, the catalogue named spectro_zcatalog_spe_quality contains all quality parameters characterising the spectra. These include the number of valid pixels (spe_npix), the minimum and maximum wavelength (spe_w_min and spe_w_max), the maximum number of individual exposures used during the combination (spe_n_dith_max), the median number of exposures calculated over the pixels of each spectrum (spe_n_dith_med), and all the flags indicating processing errors during the redshift measurement for the three classes of the classification. In Fig. 1, we show how spe_n_dith and spe_npix vary across the Euclid Deep Field North (EDF-N). The pattern resulting from the gaps between detectors is evident, and this is only partly mitigated by the Euclid dithering strategy. This clearly affects the ability to recover a redshift from spectra that fall, even in part, within these boundary areas.
@@
SPE PF employs a modified version of the Algorithm for Massive Automated Z Evaluation and Determination (AMAZED; Schmitt et al. 2019), tailored for Euclid spectroscopic properties, to calculate redshifts from spectra. This provides a broad classification of spectra into three categories: galaxies, stars, and quasars. For each category, dedicated models are fitted using a least-squares-fitting algorithm, yielding a probability distribution function (PDF) and a Bayesian evidence estimate for each category. The models for the galaxy category are a combination of six Bruzual-Charlot (BC03; Bruzual & Charlot 2003) continua, including interstellar absorption, and 13 emission-line-ratio templates built from the Virmos VLT Deep Survey (Le Fvre et al. 2013). The quasar model is a single template built from the Sloan Spectroscopic Digital Survey (SDSS) mean quasar spectrum (Vanden Berk et al. 2001), while for stars the models are a collection of 39 stellar spectra from the ESO Library of Stellar Spectra for spectrophotometric calibration (Pickles 1998). The final classification is made by selecting the category that shows the best evidence. This is collected in the spectro_zcatalog_spe_classification catalogue, together with the probability associated with each category. It should be noted that, at this stage of development and test of the spectroscopic pipeline, the quality of the classification is still being improved (see Sect. 4).
@@
For each category, the best-redshift solutions are identified by the most prominent peaks in the PDF. Up to five solutions are provided for each category, and are listed in three catalogues named spectro_zcatalog_spe_TYPE_candidates, where TYPE can be any of galaxy, star, or qso. Each object, identified by its object_id as assigned by the PF-MER pipeline, appears as many times as the number of redshift solutions, each identified by the spe_rank parameter, ranging from 0 (best solution) to at most 4. These solutions are provided for all categories, independent of the actual classification result.
@@
Figure 2: Typical spectrum (black line, uncertainties in green) of a galaxy from the EWS target sample, showing no continuum and a single emission line. From external spectroscopy, we know that this is $\mathrm{H}\,\alpha$ at $z=1.2711$; however, as discussed in the text, SPE interprets the line as [O ii] in this case and assigns an erroneous redshift.
@@ 2.2 The prior used for PDF computation
The goal of the EWS survey is to compute clustering statistics for galaxies in the range $1.880.84<z<$ 1.88, using detections of the $\mathrm{H}\,\alpha$ line at $f(\mathrm{H}\,\alpha)\,\geq 2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ to secure their redshifts. For the present paper, however, we shall work with a slightly more conservative selection specific to Q1: $1.80.9<z<$ 1.8. This constraint will be relaxed in future releases. The requirement, as defined in the original design of the survey, is that within this target range SPE PF shall deliver a sample that is 45% complete and 80% pure, for objects with an $\mathrm{H}\,\alpha$ line detected with a signal-to-noise ratio (S/N) greater than 3.5, for sources with FWHM $\leq 0.5$ arcsec (since the spectral resolution depends on the size of the object along the dispersion direction). SPE PF therefore has to supplement the measured redshifts with all the information on the measurement quality required to select a catalogue within the target range. This means supplying the correct redshift for at least 45% of the full population of $\mathrm{H}\,\alpha$ emitters matching the selection criteria, as well as a percentage of less than 20% of spurious objects with incorrect redshifts. Given the wavelength coverage of NISP and the exposure time, $\mathrm{H}\,\alpha$ is the only line detected with good S/N for a large fraction of this redshift range and for a large fraction of targeted galaxies. In addition, the continuum is barely detected, such that a galaxy at redshift $z=1.2711$ with $H_{\scriptscriptstyle\rm E}=21.5$ and an $\mathrm{H}\,\alpha$ line with flux $5\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ has a typical spectrum as shown in Fig. 2.
@@
The problem is that, when using an agnostic line identification tool, the emission line in this case is not identified as $\mathrm{H}\,\alpha$, but is incorrectly interpreted as [O ii]. In order to circumvent this issue, SPE PF introduces a prior in the calculation of the PDF, which favours solutions containing the $\mathrm{H}\,\alpha$ line. The prior acts on the PDF by dividing its value by a parameterisable factor for redshifts at which the $\mathrm{H}\,\alpha$ line can be detected. It currently has a value of 1000 below $z=0.9$ and above $z=1.8$, where $\mathrm{H}\,\alpha$ is not observed by the EWS, and unity for $1.80.9<z<$ 1.8, where it is.
@@
We thus ensure that, for objects such as the one shown in Fig. 2, the redshift solution favoured by SPE is the correct one. Of course, sometimes this will induce the opposite effect: real [O ii] emission lines can be wrongly identified as $\mathrm{H}\,\alpha$. However, to enter the red grism wavelength range, the [O ii] doublet at 3728 Å needs to come from a galaxy at $z>2.2$. At the limiting magnitude of the EWS, the number of such galaxies is lower than that of true $\mathrm{H}\,\alpha$ emitters within the target range by a factor of about 3.5 (as estimated from the EL-COSMOS data set; Saito et al. 2020) and the net result is thus an improvement of the redshift measurement success rate.

### v2

@@
The spectroscopy processing function (SPE PF) of the Euclid pipeline is dedicated to the automatic analysis of 1D spectra to determine redshifts, line fluxes, and spectral classifications. The first Euclid Quick Data Release (Q1) delivers these measurements for all $H_{\scriptscriptstyle\rm E}<22.5$ objects identified in the photometric survey. In this paper, we present an overview of the SPE PF algorithm and assess its performance by comparing its results with high-quality spectroscopic redshifts from the Dark Energy Spectroscopic Instrument (DESI) survey in the Euclid Deep Field North. Our findings highlight remarkable accuracy in successful redshift measurements, with a bias of less than $3\times 10^{-5}$ in $(z_{\rm SPE}-z_{\rm DESI})/(1+z_{\rm DESI})$ and a high precision of approximately $10^{-3}$. The majority of spectra have only a single spectral feature or none at all. To avoid spurious detections, whereby noise features are misinterpreted as lines or lines are misidentified, it is therefore essential to apply well-defined criteria on quantities such as the redshift probability, or the $\mathrm{H}\,\alpha$ flux, and signal-to-noise ratio. Using a well-tuned quality selection, we achieve an 89% redshift success rate in the target redshift range for cosmology ($0.9<z<1.8$), which is well covered by DESI for $z<1.6$. Outside this range in which the $\mathrm{H}\,\alpha$ line is observable, redshift measurements are less reliable, except for sources showing specific spectral features (e.g. two bright lines or strong continuum). The classification based on spectroscopy alone is effective for galaxies (about 80% success rate), while it is less efficient for stars and quasars ($60\%$). Ongoing refinements along the entire chain of PFs are expected to enhance both redshift measurements and spectral classification, allowing us to define the large and reliable sample required for cosmological analyses. Taking into account the planned evolution of the spectroscopic pipeline, partially based on the limitations identified in this paper, these results are encouraging for Euclid’s future galaxy clustering measurements, even though the requirements are not yet fulfilled.
@@ 1 Introduction
One of the two main objectives of the Euclid mission (Euclid Collaboration: Mellier et al., 2025) is to measure galaxy clustering in the redshift interval $0.84<z<1.88$ (hereafter, the ‘target range’), using spectroscopic redshifts measured on the spectra provided by the Near Infrared Spectro-Photometer (NISP) instrument (Euclid Collaboration: Jahnke et al., 2025). The spectra have a resolution of $R\simeq 500$, cover the wavelength range 1206–1892 $\mathrm{nm}$, and redshifts in the target range are measured by identifying the $\mathrm{H}\,\alpha$ emission line. NISP is a slitless spectrometer, which means that the spectra of all objects on the sky potentially leave a trace on the spectroscopic exposures. The only condition for a spectrum to be extracted is that a corresponding object be identified in the photometric catalogue constructed from the Euclid imaging. This paper presents spectroscopic measurements and redshifts released in the framework of the first ‘Quick Release’ of Euclid products (Euclid Quick Release Q1, 2025; Euclid Collaboration: Aussel et al., 2025), for objects brighter than $H_{\scriptscriptstyle\rm E}=22.5$ and observed over the area of the Euclid Deep Fields (EDF) in the configuration of the Euclid Wide Survey (EWS). However, given the specific objectives of the mission, in terms of redshift range and emission-line flux, as well as wavelength domain and intrinsic quality of the spectra, these data cannot be used as a homogeneous set for blind statistical studies, and special attention must be paid to redshift values outside the target range or with $\mathrm{H}\,\alpha$ lines fainter than the nominal limit of the survey. In addition, we describe the output and performance of the sole spectroscopy channel, while the final analysis of Euclid will benefit from the combination with the high-quality visible and infrared images provided by both the VIS (Euclid Collaboration: Cropper et al., 2025) and NISP instruments.
The paper is organised as follows. Section 2 describes the various spectroscopic products that are present in the Euclid scientific archive, together with the main parameters which are useful in extracting the best measurements, as well as a brief description of the methods of redshift calculation. Section 3 gives results on the efficiency and accuracy of the redshift measurements, while the results of the spectroscopic classification are presented in Sect. 4. Section 5 gives information about the quality and performance of the emission-line fluxes, and Sect. 6 discusses the quality of the redshift measurement for various sub-samples with respect to existing redshift catalogues.
@@ 2.1 SPE PF spectra handling and basic description
The Euclid data processing is handled by the Science Ground Segment, which is split into processing functions (PFs), each taking care of a specific aspect. The spectroscopy PF (SPE PF) is in charge of redshift estimation, measurement-reliability calculation, and spectroscopic classification of all objects present in the main catalogue created by the ‘merge’ PF (MER PF, Euclid Collaboration: Romelli et al., 2025). The SPE PF first reads the 1D spectra produced by the SIR PF (Euclid Collaboration: Copin et al., 2025) by combining the four individual exposures taken with different orientations of the grisms (see Euclid Collaboration: Scaramella et al., 2022 for details on the observing strategy). The data provided by the SIR PF also include flags describing the quality of each pixel in the spectrum and the number of individual exposures that actually contributed to its final value. For this Q1 release, only the spectra of objects with $H_{\scriptscriptstyle\rm E}\leq 22.5$ were extracted. At fainter magnitudes, the catalogues derived from NISP imaging suffer from significant contamination by fake sources, due to the persistence effect in infrared detectors. Spurious residuals from preceding spectroscopic exposures are detected on the following photometric images, preventing the SIR PF from reliably extracting source positions at $H_{\scriptscriptstyle\rm E}>22.5$ (see Euclid Collaboration: Polenta et al., 2025 for a description of the NIR PF and the persistence phenomenon). Before proceeding with the redshift estimate, pixels that were flagged as ‘do not use’ by the SIR PF are discarded (see Euclid Collaboration: Copin et al., 2025 for details on flags). Following tests performed on simulated data and early spectra to eliminate as many artificial features as possible, only pixels that are built from at least three exposures are retained. This is the minimum number to guarantee a proper performance of the SIR PF sigma-clipping combination.
@@
We also discard pixels that have the sixth bit of the quality flag activated: this corresponds to pixels at the edges of the spectrum, for which the absolute flux calibration curve has values close to zero, making calibration uncertain. Because of the current SIR configuration, this flag limits the available redshift interval to 0.9–1.8 in this release. Accounting also for the various edge effects (size of the fields, focal plane geometry, and gaps between detectors – see Fig. 1) yields a sample of 3.8 million processed spectra, among the 5.1 million objects in the $H_{\scriptscriptstyle\rm E}\leq 22.5$ catalogue, over the 63.1 deg2 Q1 area. This includes all spectra with at least one valid pixel. To approach a science-grade sample, a lower threshold to the number of valid pixels is clearly required; however, at this stage, we chose not to apply any such cut, since future improvements in the pipeline may lead to different values for the optimal selection to be used. The nominal length of a spectrum that would not contain any invalid pixel is about 470 pixels.
@@
It is important to note that this sample of processed (and delivered) spectra is in fact nearly two orders of magnitude larger than the number of $\mathrm{H}\,\alpha$ emitters whose redshifts are expected to be measurable at the depth of the EWS, within the Q1 area. Following the original science requirements of the mission, the NISP spectrograph was designed as to secure a redshift for an average number of $1700$ galaxies per deg2, at a flux limit of $f(\mathrm{H}\,\alpha)>2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$. Because of measuring redshifts using only the $\mathrm{H}\,\alpha$ line detection (see discussion in the following sections), by construction, these galaxies will have a redshift within the target range of $0.9<z<1.8$. In practice, this implies that, within the Q1 release, we would expect at best to recover a total of $63\times 1700\simeq 100\,000$ redshifts within the target range, which is about 2% of the total number of spectra extracted and delivered in Q1. The expected number of $\mathrm{H}\,\alpha$ emitters, in fact, should be even smaller, if we consider that the requirements are based on extracting all sources with a broad-band magnitude of $H_{\scriptscriptstyle\rm E}<24$, which is 1.5 mag deeper than the current selection. Thus, any scientific investigation based on these data requires a strong and careful selection, as we shall discuss in the following sections.
@@
To allow this selection, the catalogue named spectro_zcatalog_spe_quality contains all quality parameters characterising the spectra. These include the number of valid pixels (spe_npix), the minimum and maximum wavelength (spe_w_min and spe_w_max), the maximum number of individual exposures used during the combination (spe_n_dith_max), the median number of exposures calculated over the pixels of each spectrum (spe_n_dith_med), and all the flags indicating processing errors during the redshift measurement for the three classes of the classification. In Fig. 1, we show how spe_n_dith and spe_npix vary across the Euclid Deep Field North (EDF-N). The pattern resulting from the gaps between detectors is evident and this is only partly mitigated by the Euclid dithering strategy. An area corresponding to a lower-quality observation is apparent at about 18°$0$ and +64°30, while an additional observation, executed to correct an uncovered area, shows an excess of number of dithers and valid pixels, at about 18°20 and +65°30. This clearly affects the ability to recover a redshift from spectra that fall, even in part, within these boundary areas.
@@
The SPE PF employs a modified version of the Algorithm for Massive Automated Z Evaluation and Determination (AMAZED; Schmitt et al., 2019), tailored for Euclid spectroscopic properties, to calculate redshifts from spectra. This provides a broad classification of spectra into three categories: galaxies, stars, and quasars. For each category, dedicated models are fitted using a least-squares-fitting algorithm, yielding a probability distribution function (PDF) and a Bayesian evidence estimate for each category. The models for the galaxy category are a combination of six Bruzual–Charlot (BC03; Bruzual and Charlot, 2003) continua, including interstellar absorption, and 14 emission-line-ratio templates built from the Virmos VLT Deep Survey (VVDS, Le Fèvre et al., 2013). The initial list of BC03 contained the 39 models described in A. (2003). The 13 star formation histories included 10 instantaneous bursts at 5, 25, 100, 290, 640, 900, 1400, 2500, and 11 000 Myr, one 6-Gyr-old continuous star formation episode, and two 12-Gyr-old models with exponentially declining star formation, the time constants being $\tau=5\,{\rm Gyr}$ and $\tau=9\,{\rm Gyr}$. For each of these models, three metallicity values were used: $0.2\,Z_{\odot}$, $Z_{\odot}$, and $2.5\,Z_{\odot}$. The first test made on Euclid simulated spectra showed that, given the resolution and expected sensitivity, the metallicity diversity does not have a detectable impact on the redshift-measurement success rate. In addition, some instantaneous burst ages were very close to each other over the short wavelength range of Euclid, so we finally built the used list by keeping the six templates that were used the most and after checking that this reduced list did not induce any modification in the results. This list includes five single-burst histories at 5, 25, 640, 1400, and 5000 Myr and the exponentially declining history with $\tau=5\,{\rm Gyr}$.
The 14 emission line ratio templates cover a large spread of the ratios measured between the main emission lines in the VVDS, as well as the possible presence of fainter emission lines detected only in specific galaxy spectral types. These templates require compatible relative intensities in addition to positions of lines, and then allow fewer degrees of freedom for the fit. This prevents redshift solutions for which noise fluctuations of the spectra would be independently fitted by unrealistic emission lines. The quasar model is a single template built from the Sloan Spectroscopic Digital Survey (SDSS) mean quasar spectrum (Vanden Berk et al., 2001), while for stars the models are a collection of 36 stellar spectra from the European Southern Observatory Library of Stellar Spectra for spectrophotometric calibration (Pickles, 1998). These templates include one template for O, A, and B stars, two for F stars, two G star templates, 15 for K stars, and 12 for M stars. The cold stars are more represented because of their diversity in the spectral domain of Euclid. Ultra-cool dwarf templates (L and T) will be included in future releases. The final classification is made by selecting the category that shows the best evidence. This is collected in the spectro_zcatalog_spe_classification catalogue, together with the probability associated with each category. It should be noted that, at this stage of development and test of the spectroscopic pipeline, the quality of the classification is still being improved (see Sect. 4).
@@
For each category, the best-redshift solutions are identified by the most prominent peaks in the PDF. Up to five solutions are provided for each category and are listed in three catalogues named spectro_zcatalog_spe_TYPE_candidates, where TYPE can be any of galaxy, star, or qso. Each object, identified by its object_id as assigned by the MER PF pipeline, appears as many times as the number of redshift solutions, each identified by the spe_rank parameter, ranging from 0 (best solution) to at most 4. These solutions are provided for all categories, independent of the actual classification result. All the templates (continuum and eventually emission line ratio) that correspond to each solution are also stored and available in the Science Archive.
@@
Figure 2: Typical spectrum (black line, uncertainties in green) of a galaxy from the EWS target sample, showing no continuum and a single emission line. From external spectroscopy, we know that this is $\mathrm{H}\,\alpha$ at $z=1.2711$; however, as is discussed in the text, SPE interprets the line as [O ii] in this case and assigns an erroneous redshift.
@@ 2.2 The prior used for PDF computation
The goal of the EWS survey is to compute clustering statistics for galaxies in the range $0.84<z<1.88$, using detections of the $\mathrm{H}\,\alpha$ line at $f(\mathrm{H}\,\alpha)\,\geq 2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ to secure their redshifts. In the present paper, however, we work with a slightly more conservative selection specific to Q1: $0.9<z<1.8$. This constraint will be relaxed in future releases. The requirement, as is defined in the original design of the survey, is that, within this target range, the SPE PF will deliver a sample that is 45% complete and 80% pure, for objects with an $\mathrm{H}\,\alpha$ line detected with a signal-to-noise ratio (S/N) greater than 3.5, for sources with FWHM $\leq 0.5$ arcsec (since the spectral resolution depends on the size of the object along the dispersion direction). The SPE PF therefore has to supplement the measured redshifts with all the information on the measurement quality required to select a catalogue within the target range. This means supplying the correct redshift for at least 45% of the full population of $\mathrm{H}\,\alpha$ emitters matching the selection criteria, as well as a percentage of less than 20% of spurious objects with incorrect redshifts. Given the wavelength coverage of NISP and the exposure time, $\mathrm{H}\,\alpha$ is the only line detected with good S/N for a large fraction of this redshift range and for a large fraction of the targeted galaxies. In addition, the continuum is barely detected, such that a galaxy at redshift $z=1.2711$ with $H_{\scriptscriptstyle\rm E}=21.5$ and an $\mathrm{H}\,\alpha$ line with flux $5\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ has a typical spectrum as the one shown in Fig. 2.
@@
The problem is that, when using an agnostic line identification tool, the emission line in this case is not identified as $\mathrm{H}\,\alpha$, but is incorrectly interpreted as [O ii]. In order to circumvent this issue, SPE PF introduces a prior in the calculation of the PDF, which favours solutions containing the $\mathrm{H}\,\alpha$ line. The prior acts on the PDF by dividing its value by a tunable factor for redshifts at which the $\mathrm{H}\,\alpha$ line can be detected. It currently has a value of 1000 below $z=0.9$ and above $z=1.8$, where $\mathrm{H}\,\alpha$ is not observed by the EWS, and unity for $0.9<z<1.8$, where it is.
@@
We thus ensure that, for objects such as the one shown in Fig. 2, the redshift solution favoured by SPE is the correct one. Of course, sometimes this will induce the opposite effect: real [O ii] emission lines can be wrongly identified as $\mathrm{H}\,\alpha$. However, to enter the red grism wavelength range, the [O ii] doublet at 3728 Å needs to come from a galaxy at $z>2.2$. At the limiting magnitude of the EWS, the number of such galaxies is lower than that of true $\mathrm{H}\,\alpha$ emitters within the target range by a factor of about 3.5 (as estimated from the EL-COSMOS dataset; Saito et al., 2020) and the net result is thus an improvement of the redshift measurement success rate.